# On the dimension distortion under fractionally smooth mappings

### v1

@@ 2.2. Mappings between metric spaces
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two metric spaces. Given $\alpha>0$, a mapping $f:X\rightarrow Y$ and a set $B\subset X$, we define the $\alpha$-Hölder coefficient of $f$ on $B$ as
@@
Given an at most countable index set $I$, we denote by $\ell^{p}(I)$ the space of real-valued sequences $\{c_{i}\}_{i\in I}$ with finite $p$-norm $(\sum_{i\in I}c_{i}^{p})^{1/p}<\infty$. We call $\sum_{i\in I}c_{i}^{p}$ the $p$-sum of the sequence $\{c_{i}\}_{i\in\mathbb{N}}$.
@@
Let $f:X\rightarrow Y$ be a mapping between two arbitrary metric spaces. For $p\in(1,\infty)$ and $\alpha>0$, we say $f$ is $(p,\alpha)$-compactly Hölder, and write $f\in CH^{p,\alpha}(X:Y)$, if for any compact set $E\subset X$ and any $\varepsilon\in(0,1)$ there are $r_{E}>0$ and $C_{E}>0$ satisfying the following:
if $\{B_{i}\}_{i\in I}$ is a collection of balls $B_{i}:=B(x_{i},r)$ with $x_{i}\in X$, $r<r_{E}$ that covers $E$ and $B(x_{i},\varepsilon r)\cap B(x_{j},\varepsilon r)=\emptyset$ for all distinct $i,j\in I$, then the $p$-sum of the Hölder coefficients of $f$ on $B_{i}$ is at most $C_{E}$, i.e.,
@@
Here we follow the convention that if $\{B(x_{i},r)\}_{i\in I}$ covers $E$, it is implied that $B(x_{i},r)\cap E\neq\emptyset$ for all $i$, but not all $x_{i}$ necessarily lie in $E$. Note that applying the definition on singleton sets yields that compactly Hölder mappings are Hölder continuous on compact sets. Moreover, in the setting of Definition 2.4 it is actually implied by (2.9) that there are $C_{i}>0$ such that
(2.10) $|f(B(x_{i},r)|\leq C_{i}|B(x_{i},r)|^{\alpha}$
with $\sum_{i\in I}C_{i}^{p}\leq C_{E}$. This inequality provides insight on the relation with the Euclidean setting. More specifically, it hints at how the motivation for Definition 2.4 comes from continuous super-critical Sobolev maps between Euclidean spaces, i.e. continuous maps in $W^{1,p}(\Omega;\mathbb{R}^{n})$ with $\Omega\subset\mathbb{R}^{n}$ and $p>n$. For more details on the motivation behind these mappings see [21].
@@
Given $f\in CH^{p,\alpha}(X:Y)$, a non-empty compact set $E\subset X$, some $\varepsilon\in(0,1)$ and a cover of $E$ by balls $\{B(x_{i},r_{i})\}_{i\in I}$ with $r_{i}\leq r_{E}$ and $B(x_{i},\varepsilon r_{i})\cap B(x_{j},\varepsilon r_{j})=\emptyset$ for all distinct $i,j\in I$, then there are $C_{i}>0$ with
(2.11) $|f(B(x_{i},r_{i})|\leq C_{i}|B(x_{i},r_{i})|^{\alpha}$
and $\sum_{i\in I}C_{i}^{p}\leq C_{E}$. This change from radii $r$ to not necessarily equal radii $r_{i}$ can be made by bounding the partial sum $\sum_{i\in\{i_{1},\dots i_{n}\}}|f|_{\alpha,B(x_{i},r_{i})}$ by a sum of the $\alpha$-Hölder semi-norms over balls $B(y_{j},\rho)$, $\rho=\min\{r_{i}:i\in\{i_{1},\dots i_{n}\}\}$, which cover $\cup_{i\in\{i_{1},\dots i_{n}\}}B(x_{i},r_{i})$ and are $\varepsilon$-disjoint, with the latter sum being by definition at most $C_{E}$, for all $n\in\mathbb{N}$.
@@
for every $x,y\in X\setminus E$. The collection of all the $s$-gradients of $u$ is denoted by $\mathcal{D}^{s}(u)$. Given $p\in(0,\infty)$, the (homogeneous) fractional Hajłasz–Sobolev space $\dot{M}^{s,p}(X:Y)$ is defined as the collection of all the measurable functions $u\colon X\rightarrow Y$ such that
$\|u\|_{\dot{M}^{s,p}(X:Y)}:=\inf_{g\in\mathcal{D}^{s}(u)}\|g\|_{L^{p}(X)}<\infty.$
Here and thereafter, we make the agreement that $\inf\emptyset:=\infty$.
@@
A few comments are in order. We use the term ‘semi-norms’ for $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$, $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$, and $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$ even though the triangle inequality holds only when $p,q\geq 1$. A genuine ‘norm’ can be obtained by passing to the quotient space modulo constant functions. Since altering functions on sets of measure zero do not affect their membership in $\dot{M}^{s,p}(X:Y)$, $\dot{M}^{s}_{p,q}(X:Y)$, or $\dot{N}^{s}_{p,q}(X:Y)$, it is standard to regard these spaces as consisting of equivalence classes of functions. We adopt this convention here, but we choose to omit the details.
When $Y=\mathbb{R}$, we make the following abbreviations: $\dot{M}^{s,p}(X):=\dot{M}^{s,p}(X:\mathbb{R})$ $\dot{M}^{s}_{p,q}(X):=\dot{M}^{s}_{p,q}(X:\mathbb{R})$, and $\dot{N}^{s}_{p,q}(X):=\dot{N}^{s}_{p,q}(X:\mathbb{R})$. It was shown in [57] that $\dot{M}^{s}_{p,q}(\mathbb{R}^{n})$ coincides with the classical Triebel–Lizorkin space $\dot{F}^{s}_{p,q}(\mathbb{R}^{n})$ for any $s\in(0,1)$, $p\in(\frac{n}{n+s},\infty)$, and $q\in(\frac{n}{n+s},\infty]$, and $\dot{N}^{s}_{p,q}(\mathbb{R}^{n})$ coincides with the classical Besov space $\dot{B}^{s}_{p,q}(\mathbb{R}^{n})$ for any $s\in(0,1)$, $p\in(\frac{n}{n+s},\infty)$, and $q\in(0,\infty]$. In particular, the Hajłasz–Triebel–Lizorkin and Hajłasz–Besov spaces on $\mathbb{R}^{n}$ contain the classical fractional Sobolev spaces as special cases (see [37, Chapter 2]). When $s=1$, we have that $\dot{M}^{1}_{p,\infty}(\mathbb{R}^{n})=\dot{M}^{1,p}(\mathbb{R}^{n})$ coincides with the classical Sobolev space $\dot{W}^{1,p}(\mathbb{R}^{n})$ for any $p\in(1,\infty)$; see Lemma 4.5 and [39]. We refer the reader to [38, 39, 64, 57, 43, 3, 42, 5, 4] for more information on Sobolev, Triebel–Lizorkin, and Besov spaces on metric measure spaces.
The following notions for measures are also typically needed in this setting. We say a metric measure space $(X,d,\mu)$ is locally $Q$-homogeneous, for some $Q>0$, if for all compact $K\subset X$ there are constants $\tilde{R}_{\text{hom}}(K)>0$, $\tilde{C}_{\text{hom}}(K)\geq 1$ such that
@@
whenever $B(x,r)$ is a ball centered at a point in $K$ with $r\leq R$. We say a metric space $(X,d)$ is locally $Q$-homogeneous if there is a measure $\mu$ on $X$ such that $(X,d,\mu)$ is locally $Q$-homogeneous. One particular property due to local homogeneity that we need is the lower bound on the measure of a ball by its radius to a power. More specifically, for $R_{\text{hom}}(K)=\tilde{R}_{\text{hom}}(K)/3$ and a potentially larger $C_{\text{hom}}(K)$, it can be shown that
@@
It can be necessary at times to also have a similar upper bound on the measure. We say that $(X,d,\mu)$ is $Q$-Ahlfors regular for some $Q>0$ if there is a constant $C_{A}>0$ such that for all $x\in X$ and all $r\in(0,|X|)$ we have
@@ 3. Intermediate dimensions under compactly Hölder mappings
If $Y$ is uniformly perfect with constant $c_{u}^{\prime}>0$, then for every $F\subset Y$ and every $M>0$ with $|F|<M<Mc_{u}^{\prime-1}<|Y|$, there is a set $L(F)$ containing $F$ with
@@
Since all mappings considered henceforth are defined on $X$ and into $Y$, we set $CH^{p,\alpha}=CH^{p,\alpha}(X:Y)$. Let $f\in CH^{p,\alpha}$ for $p\in(1,\infty)$, $\alpha>0$ and $E\subset X$ be a bounded set. We plan on using coverings of $E$ by dyadic cubes, provided by Proposition 2.2, and determine admissible coverings for $\dim_{\theta}f(E)$ based on those of $E$ and shrinking properties of $f$ controlled by inequality (2.11). Before we delve into the proof, we give an outline of the combinatorial part of our method. The main idea is to fix an arbitrary $d>\dim_{\theta}E$, and for every $\varepsilon>0$ fix for every $\delta\in(0,\delta_{\varepsilon})$ a dyadic covering $\mathcal{Q}_{\delta}$ of $E$ as in Proposition 2.2. The images of the covering cubes form a cover of $f(E)$. After determining a $\delta_{Y}$ to use for $\dim_{\theta}f(E)$ (see (3.3)), we need to modify the cubes in order to get a $\delta_{Y}^{1/\theta}$-admissible cover of $f(E)$. To do so, we build a combinatorial graph with cubes as vertices, in the following manner (see Figure 1):
@@
Based on the above graph, the strategy then is to use a covering of $f(E)$ consisting of images of all blue cubes, and “enlarged” images of green cubes (using Lemma 3.1). This by construction is a $\delta_{Y}^{1/\theta}$-admissible covering of $f(E)$, with the least cardinality possible based on $\mathcal{Q}_{\delta}$, and with a combinatorial representation that facilitates estimates on the sum of respective diameters to an appropriate exponent, achieving the desired upper bound on $\dim_{\theta}f(E)$. With this strategy in mind, we are ready to delve into the proof of Theorem 1.1.
@@
We focus on the case $d_{E}(\theta)\leq pd_{E}(\theta)(\alpha p+d_{E}(\theta))^{-1}$, and the other case can be treated similarly. Suppose $\dim_{\theta}E=d_{E}<d<d^{\prime}<p-\alpha p$, and set $D=\tfrac{p{d^{\prime}}}{\alpha p+d^{\prime}}$. The choice $d<d^{\prime}<p-\alpha p$ guarantees that $d\neq D$, but the proof is similar in the case $d_{E}\geq p-\alpha p$, by choosing $d>d_{E}$.
@@
We say a cube $Q$ that intersects $E$ and is contained in some $Q_{i}^{k_{i}}\in\mathcal{Q}_{\delta}$, for some $i\in I_{\delta}$, is blue if $|f(Q)|\in[\delta_{Y}^{1/\theta},\delta_{Y})$, green if $|f(Q)|<\delta_{Y}^{1/\theta}$, and red otherwise. Notice that the terminology applies not only to cubes in $\mathcal{Q_{\delta}}$, but also to their sub-cubes. In addition, it applies only to the aforementioned types of cubes, due to our initial reduction on $\mathcal{Q}_{\delta}$ and by Theorem A (i). By uniform continuity of $f$, it is guaranteed that past a certain level, all sub-cubes of cubes in $\mathcal{Q_{\delta}}$ will be blue or green. The strategy is to sub-divide every red cube in $\mathcal{Q_{\delta}}$ and all their red sub-cubes the least number of times necessary, in order for the resulting cubes to all be blue or green. Then, apply Lemma 3.1 to the images of all green cubes under $f$, in order to “enlarge” them, and pick a covering of $f(E)$ consisting of these enlarged sets and the images of all blue cubes under $f$. By construction, this covering is a $\delta_{Y}^{1/\theta}$-admissible covering for $\dim_{\theta}f(E)$. Then, it would be enough to show that the sum of their diameters is small enough.
Suppose $m_{K_{\delta}}\geq m_{N_{\delta}}$ is the smallest integer such that all cubes of level $m_{K_{\delta}}$ which intersect $E$ and are contained in $\bigcup_{i\in I_{\delta}}Q_{i}^{k_{i}}$, are either green or blue. Note that $m_{K_{\delta}}$ could be strictly larger than $m_{N_{\delta}}$, because we can have red $m_{N_{\delta}}$-level cubes in $\mathcal{Q_{\delta}}$, or red cubes in $\mathcal{Q_{\delta}}$ might need to be subdivided past the level $m_{N_{\delta}}$ to give only blue and green descendants. Based on this subdivision of red cubes in $\mathcal{Q_{\delta}}$ and their red sub-cubes, solely dependent on the way $f$ distorts diameters, we build a directed graph $\mathcal{G}=\mathcal{G}(\mathcal{Q_{\delta}})$. For an integer $m\in[m_{1},m_{K_{\delta}}]$, set $V_{m}$ to be the collection of all $m$-level cubes intersecting $E$ and contained in $\cup_{i\in I_{\delta}}Q_{i}^{k_{i}}$ that are either red, or are contained in a red cube of level m1m-1. Then, the vertex set of $\mathcal{G}$ is defined to be the collection of cubes
@@
Note that blue and green cubes in $\mathcal{Q_{\delta}}$ are part of no edge, because there is no need for them to be sub-divided, and they cannot have an ancestor also in $\mathcal{Q}_{\delta}$ by the respective reduction in the beginning of the proof. Let $V_{R},V_{B}^{\text{orig}},V_{B}^{\prime},V_{G}$ denote the red cubes in $V$, blue cubes in $V\cap\mathcal{Q_{\delta}}$, blue cubes in $V$ that do not lie in $\mathcal{Q_{\delta}}$ (i.e. they have an ancestor in $V$), and green cubes in $V$, respectively. Notice that if we have a bound on $\text{card}V_{R}$, we can also bound $\text{card}V_{B}^{\prime}$ by Remark 2.1:
@@
because every vertex in $V_{B}^{\prime}$ is connected to a red vertex in $V$, and for each red $Q\in V$ there are at most $N_{d}$ cubes $\tilde{Q}\in V$ with $(Q,\tilde{Q})\in E$.
@@
where $c_{u,0}\varepsilon\delta^{-d/\theta}$ bounds the number of green cubes in $\mathcal{Q}_{\delta}$, and $N_{d}\,\text{card}V_{R}$ bounds the number of green cubes in $V$ which are connected to some red vertex in $V$ by an edge, after another application of Remark 2.1. Let
@@
Let $\{Q_{j}^{m}\}_{j\in J_{m}}$ be all red $m$-level cubes. If $Q_{j}^{m}$ is such a cube, by $f\in CH^{p,\alpha}$ and (2.11) we have
@@
By choice of $\delta_{Y}$ in (3.3), $\delta_{Y}=\delta^{d/D}$, we have
@@
    Comparing our terminology to that of Kaufman, red and blue cubes correspond to “major” and “minor” cubes in [54]. Due to the nature of the intermediate dimension, requiring a lower bound on the diameters of covering sets, we had to introduce the class of green cubes to account for sets in the target that are not too large, but are too small. This is not necessary when investigating the distortion of the upper box-counting dimension, and thus a similar notion was not needed in [54].

### v2

@@ 2.2. Mappings between metric spaces
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two metric spaces. Given $\alpha\in(0,\infty)$, a mapping $f:X\rightarrow Y$ and a set $B\subset X$, we define the $\alpha$-Hölder coefficient of $f$ on $B$ as
@@
Given an at most countable index set $I$, we denote by $\ell^{p}(I)$, $p\in(0,\infty)$, the space of real-valued sequences $\{c_{i}\}_{i\in I}$ with finite $p$-norm $(\sum_{i\in I}c_{i}^{p})^{1/p}<\infty$. We call $\sum_{i\in I}c_{i}^{p}$ the $p$-sum of the sequence $\{c_{i}\}_{i\in\mathbb{N}}$.
@@
Let $f:X\rightarrow Y$ be a mapping between two arbitrary metric spaces. For $p,\alpha\in(0,\infty)$, we say $f$ is $(p,\alpha)$-compactly Hölder, and write $f\in CH^{p,\alpha}(X:Y)$, if for any compact set $E\subset X$ and any $\varepsilon\in(0,1)$ there are $r_{E}>0$ and $C_{E}>0$ satisfying the following:
if $\{B_{i}\}_{i\in I}$ is a collection of balls $B_{i}:=B(x_{i},r_{i})$ with $x_{i}\in X$, $r_{i}<r_{E}$ that covers $E$ and $B(x_{i},\varepsilon r_{i})\cap B(x_{j},\varepsilon r_{j})=\emptyset$ for all distinct $i,j\in I$, then the $p$-sum of the Hölder coefficients of $f$ on $B_{i}$ is at most $C_{E}$, i.e.,
@@
Here we follow the convention that if $\{B(x_{i},r_{i})\}_{i\in I}$ covers $E$, it is implied that $B(x_{i},r_{i})\cap E\neq\emptyset$ for all $i$, but not all $x_{i}$ necessarily lie in $E$. Note that applying the definition on singleton sets yields that compactly Hölder mappings are Hölder continuous on compact sets. Moreover, in the setting of Definition 2.4 it is actually implied by (2.9) that there are $C_{i}>0$ such that
(2.10) $|f(B(x_{i},r_{i})|\leq C_{i}|B(x_{i},r_{i})|^{\alpha}$
with $\sum_{i\in I}C_{i}^{p}\leq C_{E}$. This inequality provides insight on the relation with the Euclidean setting. More specifically, it hints at how the motivation for Definition 2.4 comes from continuous super-critical Sobolev maps between Euclidean spaces, i.e. continuous maps in $W^{1,p}(\Omega;\mathbb{R}^{n})$ with $\Omega\subset\mathbb{R}^{n}$ and $p>n$. See also [22].
@@
In [22] the class of compactly Hölder maps was initially defined with $\{B_{i}\}$ being a cover where all balls are of the same radius $r$ (see [22, Definition 2.3]). While that condition is enough to determine the distortion of the Minkowski dimension, all types of Sobolev maps considered in this manuscript and in [22] satisfy the stronger condition with potentially distinct radii $r_{i}$. Thus, we decided to update the definition to require this stronger property, especially since the motivation is to appropriately generalize these mapping classes to a solely metric setting (with no measure). It should be noted that all results in [22] stated for the “older” and wider class of compactly-Hölder mappings also hold for the notion in Definition 2.4.
@@
for every $x,y\in X\setminus E$. The collection of all the $s$-gradients of $u$ is denoted by $D^{s}(u)$. Given $p\in(0,\infty)$, the (homogeneous) fractional Hajłasz–Sobolev space $\dot{M}^{s,p}(X:Y)$ is defined as the collection of all the measurable functions $u\colon X\rightarrow Y$ such that
$\|u\|_{\dot{M}^{s,p}(X:Y)}:=\inf_{g\in D^{s}(u)}\|g\|_{L^{p}(X)}<\infty.$
Here and thereafter, we make the agreement that for semi-norms $\inf\emptyset:=\infty$.
@@
A few comments are in order. We use the term ‘semi-norms’ for $\|u\|_{\dot{M}^{s,p}(X:Y)}$, $\|\cdot\|_{\dot{M}^{s}_{p,q}(X:Y)}$, and $\|u\|_{\dot{N}^{s}_{p,q}(X:Y)}$ even though the triangle inequality holds only when $p,q\geq 1$. A genuine ‘norm’ can be obtained by passing to the quotient space modulo constant functions. Since altering functions on sets of measure zero do not affect their membership in $\dot{M}^{s,p}(X:Y)$, $\dot{M}^{s}_{p,q}(X:Y)$, or $\dot{N}^{s}_{p,q}(X:Y)$, it is standard to regard these spaces as consisting of equivalence classes of functions. We adopt this convention here, but we choose to omit the details.
When $Y=\mathbb{R}$, we make the following abbreviations: $\dot{M}^{s,p}(X):=\dot{M}^{s,p}(X:\mathbb{R})$ $\dot{M}^{s}_{p,q}(X):=\dot{M}^{s}_{p,q}(X:\mathbb{R})$, and $\dot{N}^{s}_{p,q}(X):=\dot{N}^{s}_{p,q}(X:\mathbb{R})$. It was shown in [59] that $\dot{M}^{s}_{p,q}(\mathbb{R}^{n})$ coincides with the classical Triebel–Lizorkin space $\dot{F}^{s}_{p,q}(\mathbb{R}^{n})$ for any $s\in(0,1)$, $p\in(\frac{n}{n+s},\infty)$, and $q\in(\frac{n}{n+s},\infty]$, and $\dot{N}^{s}_{p,q}(\mathbb{R}^{n})$ coincides with the classical Besov space $\dot{B}^{s}_{p,q}(\mathbb{R}^{n})$ for any $s\in(0,1)$, $p\in(\frac{n}{n+s},\infty)$, and $q\in(0,\infty]$. In particular, the Hajłasz–Triebel–Lizorkin and Hajłasz–Besov spaces on $\mathbb{R}^{n}$ contain the classical fractional Sobolev spaces as special cases (see [39, Chapter 2]). When $s=1$, we have that $\dot{M}^{1}_{p,\infty}(\mathbb{R}^{n})=\dot{M}^{1,p}(\mathbb{R}^{n})$ coincides with the classical Sobolev space $\dot{W}^{1,p}(\mathbb{R}^{n})$ for any $p\in(1,\infty)$; see Lemma 4.5 and [41]. We refer the reader to [40, 41, 66, 59, 45, 3, 44, 5, 4] for more information on Sobolev, Triebel–Lizorkin, and Besov spaces on metric measure spaces.
The following notions for measures are also typically assumed in this setting. We say a metric measure space $(X,d,\mu)$ is locally $Q$-homogeneous, for some $Q>0$, if for all compact $K\subset X$ there are constants $\tilde{R}_{\text{hom}}(K)>0$, $\tilde{C}_{\text{hom}}(K)\geq 1$ such that
@@
whenever $B(x,r)$ is a ball centered at a point in $K$ with $r\leq\tilde{R}_{\text{hom}}(K)$. We say a metric space $(X,d)$ is locally $Q$-homogeneous if there is a measure $\mu$ on $X$ such that $(X,d,\mu)$ is locally $Q$-homogeneous. One particular property due to local homogeneity that we need is the lower bound on the measure of a ball by its radius to a power. More specifically, for $R_{\text{hom}}(K)=\tilde{R}_{\text{hom}}(K)/3$ and a potentially larger $C_{\text{hom}}(K)$, it can be shown that
@@
It can be necessary at times to also have a similar upper bound on the measure, and especially in the study of quasisymmetric mappings. We say that $(X,d,\mu)$ is $Q$-Ahlfors regular for some $Q>0$ if there is a constant $C_{A}>0$ such that for all $x\in X$ and all $r\in(0,|X|)$ we have
@@ 3. Intermediate dimensions under compactly Hölder mappings
If $Y$ is $c_{u}^{\prime}$-uniformly perfect, then for every $F\subset Y$ and every $M>0$ with $|F|<M<Mc_{u}^{\prime-1}<|Y|$, there is a set $L(F)$ containing $F$ with
@@
Since all mappings considered henceforth are defined on $X$ and into $Y$, we set $CH^{p,\alpha}=CH^{p,\alpha}(X:Y)$. Let $f\in CH^{p,\alpha}$ for $p\in(1,\infty)$, $\alpha>0$ and $E\subset X$ be a bounded set. We plan on using coverings of $E$ by dyadic cubes, provided by Proposition 2.2, and determine admissible coverings for $\dim_{\theta}f(E)$ based on those of $E$ and shrinking properties of $f$ controlled by inequality (2.10). Before we delve into the proof, we give an outline of the combinatorial part of our method. The main idea is to fix an arbitrary $d>\dim_{\theta}E$, and for every $\varepsilon>0$ fix for every $\delta\in(0,\delta_{\varepsilon})$ a dyadic covering $\mathcal{Q}_{\delta}$ of $E$ as in Proposition 2.2. The images of the covering cubes form a cover of $f(E)$. After determining a $\delta_{Y}$ to use for $\dim_{\theta}f(E)$ (see (3.3)), we need to modify the cubes in order for their images to yield a $\delta_{Y}^{1/\theta}$-admissible cover of $f(E)$. To do so, we build a combinatorial graph with cubes as vertices, in the following manner (see Figure 1):
@@
Based on the above graph, the strategy then is to use a covering of $f(E)$ consisting of images of all blue cubes, and “enlarged” images of green cubes (using Lemma 3.1). This by construction is a $\delta_{Y}^{1/\theta}$-admissible covering of $f(E)$, with the least cardinality possible among coverings based on $\mathcal{Q}_{\delta}$, and with a combinatorial representation that facilitates estimates on the sum of respective diameters to an appropriate exponent, achieving the desired upper bound on $\dim_{\theta}f(E)$. With this strategy in mind, we are ready to delve into the proof of Theorem 1.1.
@@
Fix $\theta\in(0,1)$. We focus on the case $\dim_{\theta}E=d_{E}(\theta)<pd_{E}(\theta)(\alpha p+d_{E}(\theta))^{-1}$, and the other case can be treated similarly. Suppose $d_{E}:=d_{E}(\theta)<d<d^{\prime}<p-\alpha p$, and set $D=\tfrac{p{d^{\prime}}}{\alpha p+d^{\prime}}$. The choice $d<d^{\prime}<p-\alpha p$ guarantees that $d\neq D$, but the proof is similar in the case $d_{E}(\theta)\geq pd_{E}(\theta)(\alpha p+d_{E}(\theta))^{-1}$, by choosing arbitrary $D>d^{\prime}>d>d_{E}$.
@@
We say a cube $Q$ that intersects $E$ and is contained in some $Q_{i}^{k_{i}}\in\mathcal{Q}_{\delta}$, for some $i\in I_{\delta}$, is blue if $|f(Q)|\in[\delta_{Y}^{1/\theta},\delta_{Y})$, green if $|f(Q)|<\delta_{Y}^{1/\theta}$, and red otherwise. Notice that the terminology applies not only to cubes in $\mathcal{Q_{\delta}}$, but also to their sub-cubes. In addition, it applies only to the aforementioned types of cubes, due to our initial reduction on $\mathcal{Q}_{\delta}$ and by Theorem A (i). By uniform continuity of $f$, it is guaranteed that past a certain level, all sub-cubes of cubes in $\mathcal{Q_{\delta}}$ will be blue or green. The strategy is to sub-divide every red cube in $\mathcal{Q_{\delta}}$ and all their red sub-cubes the least number of times necessary, in order for the resulting cubes that intersect $E$ to all be blue or green. Then, apply Lemma 3.1 to the images of all green cubes under $f$, in order to “enlarge” them, and pick a covering of $f(E)$ consisting of these enlarged sets and the images of all blue cubes under $f$. By construction, this covering is a $\delta_{Y}^{1/\theta}$-admissible covering for $\dim_{\theta}f(E)$. Then, it would be enough to show that the sum of their diameters raised to $D$ is small enough.
Suppose $m_{K_{\delta}}\geq m_{N_{\delta}}$ is the smallest integer such that all cubes of level $m_{K_{\delta}}$ which intersect $E$ and are contained in $\bigcup_{i\in I_{\delta}}Q_{i}^{k_{i}}$, are either green or blue. Note that $m_{K_{\delta}}$ could be strictly larger than $m_{N_{\delta}}$, because we can have red $m_{N_{\delta}}$-level cubes in $\mathcal{Q_{\delta}}$, or red cubes in $\mathcal{Q_{\delta}}$ might need to be subdivided past the level $m_{N_{\delta}}$ to give only blue and green descendants. Based on this subdivision of red cubes in $\mathcal{Q_{\delta}}$ and their red sub-cubes, solely dependent on the way $f$ distorts diameters, we build a directed graph $\mathcal{G}=\mathcal{G}(\mathcal{Q_{\delta}})$. For an integer $m\in[m_{1},m_{K_{\delta}}]$, set $V_{m}$ to be the collection of all $m$-level cubes intersecting $E$ and contained in $\cup_{i\in I_{\delta}}Q_{i}^{k_{i}}$ that are either red, or are contained in a red cube of level m1m-1. This excludes descendants of blue and green cubes from being included in $V_{m}$, because those are not subdivided. Then, the vertex set of $\mathcal{G}$ is defined to be the collection of cubes
@@
Note that blue and green cubes in $\mathcal{Q_{\delta}}$ are part of no edge, because there is no need for them to be sub-divided, and they cannot have an ancestor also in $\mathcal{Q}_{\delta}$ by the respective reduction in the beginning of the proof. Let $V_{R},V_{B}^{\text{orig}},V_{B}^{\prime},V_{G}$ denote the red cubes in $V$, blue cubes in $V\cap\mathcal{Q_{\delta}}$, blue cubes in $V$ that do not lie in $\mathcal{Q_{\delta}}$ (i.e. blue cubes with an ancestor in $V$), and green cubes in $V$, respectively. Notice that if we have a bound on $\text{card}V_{R}$, we can also bound $\text{card}V_{B}^{\prime}$ by Remark 2.1:
@@
because every vertex in $V_{B}^{\prime}$ is connected to a red ancestor vertex in $V$, and for each red $Q\in V$ there are at most $N_{d}$ cubes $\tilde{Q}\in V$ with $(Q,\tilde{Q})\in E$.
@@
where $c_{u,0}\varepsilon\delta^{-d/\theta}$ bounds the number of green cubes in $\mathcal{Q}_{\delta}$, and $N_{d}\,\text{card}V_{R}$ bounds the number of green cubes in $V$ which are connected to some red ancestor vertex in $V$ by an edge, after another application of Remark 2.1. Let
@@
Let $\{Q_{j}^{m}\}_{j\in J_{m}}$ be all red $m$-level cubes in $V$. If $Q_{j}^{m}$ is such a cube, by $f\in CH^{p,\alpha}$ and (2.10) we have
@@
By choice of $\delta_{Y}=\delta^{d/D}$ in (3.3), we have
@@
    Comparing our terminology to that of Kaufman, red and blue cubes correspond to “major” and “minor” cubes in [56]. Due to the nature of the intermediate dimension, requiring a lower bound on the diameters of covering sets, we had to introduce the class of green cubes to account for sets in the target that are not too large, but are too small. This is not necessary when investigating the distortion of the upper box-counting dimension, and thus a similar notion was not needed in [56]. We also note that “critical” in [56] corresponds to “red cubes with blue or green descendants” in our language, which is, however, not a necessary notion for our argument.